# Fairness And Performance In Harmony: Data Debiasing Is All You Need

### v1

@@ VII Conclusion
Fairness in decision-making is critical, whether conducted by machine learning models or human experts. Our work first investigates individual fairness, using a consistency score to evaluate decision-making among a diverse group of subjects. Experimental results reveals that ML models such as XGBoost and BiLSTM achieve significantly higher consistency scores than human experts. This finding suggests that ML models can provide more reliable and fair treatment of similar applications, addressing the subjectivity and cognitive biases inherent in human decision-making.
In examining group fairness, we propose a gender-debiasing pipeline that effectively removes gender-specific language from the features without compromising prediction performance. Our experimental results demonstrate that all three models such as XGBoost, BiLSTM and KNN maintain or improve on classification accuracy using debiased data. This validates our hypothesis that fairness and performance can coexist, challenging the assumed trade-off between the two.
@@
Our findings have practical implications for the design of admission systems, suggesting that ML models could serve as valuable decision support tools, particularly in initial screening stages where consistency is crucial. The models could help identify potential inconsistencies in human decisions, prompting additional review in cases where machine and human assessments significantly diverge. Furthermore, the quantification of consistency scores provides a new dimension for evaluating and improving admission processes, beyond traditional accuracy metrics.
Moving forward, we could extend this work by examining biases from the examiners’ perspective and other demographic traits such as nationality, race, and religion. Furthermore, we could investigate potential bias introduced by PLM used for text embedding. Finally, we could explore leveraging Large Language Models (LLMs) for debiasing data to enhance the efficiency and accuracy of the debiasing process.

### v2

@@ 5 Conclusion
Fairness in decision-making is critical, whether conducted by artificial intelligence or human. Our work investigates individual fairness, using a consistency score to evaluate decision-making among a diverse group of subjects. Experimental results reveals that ML models such as XGBoost and BiLSTM achieve significantly higher consistency scores than human experts. This finding suggests that ML models can provide more consistent treatment of similar applications, addressing any potential subjectivity and cognitive biases inherent in human decision-making.
@@
Our findings have practical implications for the design of admission systems, suggesting that ML models could serve as valuable decision support tools, particularly in initial screening stages where consistency is crucial. The models could help identify potential inconsistencies in human decisions, prompting additional review in cases where machine and human assessments diverge. Furthermore, the quantification of consistency scores provides a new dimension for evaluating and improving admission processes, beyond traditional accuracy metrics.
Acknowledgments. This research is supported in part by the Ministry of Education, Singapore (MOE), under its Academic Research Fund Tier 2 (Award No. MOE-T2EP20123-0015), and the Singapore University of Technology and Design (SUTD) under grant RS-MEFAI-00011. Any opinions, findings and conclusions, or recommendations expressed in this material are those of the authors and do not reflect the views of MOE or SUTD.